# Rui Eduardo Lopes, Student Member, IEEE, Susana Sargento, Senior Member, IEEE

### v1

@@ VI-A Overall end-to-end interaction
The system to be researched orchestrate the service deployment, execution and evaluation of intra- and inter- vehicular services. Services can be associated with domains, that define not only requirements but the universe of information the service can subscribe. Service developers define a service descriptor (an example is detailed in 1) to describe the main service characteristics. This set of requirements is used by the orchestrator to configure and dispatch the services to the appropriated hypervisor in the appropriated ECUs and configure the datapath, i.e. the network and data layer resources required for the service to communicate with others.
@@
Services must be deployed by the orchestrator which must decide the best node(s) for the service characteristics — automatically, manually, or based on criteria (e.g. specifications, time, energy, cost), and ensure latency bounds for the deployment of the service and for the exchange of messages between them if the service descriptor (potential example on 1) requires criticality and/or QoS levels on the data exchange. The orchestrator is the module responsible for a) evaluating the nodes on the match between the service characteristics and the node characteristics, and b) dispatching the configurations to the $N$ nodes best fulfilling the service characteristics. Receiving as input both the service descriptor and the node characteristics, the algorithm outputs a score to each node and if the node may be used or not (to support requirements that must be fulfilled).
@@
The vehicular service developer shall interact with the system through a declarative paradigm, i.e. set the wanted state [^62] Kubernetes is a widely known example of such declarative paradigm.. A file — or equivalent structure — describing the necessary characteristics of the vehicular service can support this paradigm. This information is defined by the service developer and can be used by the system to be developed to support the service deployment.
@@
Associated with each field is not only the value, with the type prescribed, but also its criticality. Since not all conditions can be met by a node, it is essential for the orchestrator module to know which metrics can be unmet and by what margin.
@@
For example, when considering the use case in section V-B, one can imagine the service developer creating a service descriptor such as the one in 1. This service descriptor is then submitted to a platform that sends to the controller within the target vehicle(s). Within each vehicle a controller is responsible fo evaluating the ECUs and configuring them and the necessary network components to ensure that the service requirements are properly considered. In the case of the considered service, this entails at least deploying the necessary containers within an ECU connected to the necessary actuators - the motors for the fixing points and the ramp - and configuring the switches connecting such actuators and the ECUs where the service connect with the necessary TSN TAS scheduling - while keeping fulfilling the other already running services and their requirements.
@@
VI-B Proposal of evaluation architecture
The evaluation of the proposed system requires at least a testbed with a topology equivalent to the one depicted in fig. 22. A smart city infrastructure, including its cloud and RSUs form a Deterministic Network domain. This domain is responsible for ensuring a deterministic end-to-end latency for messages of vehicular servics by configuring appropriately the necessary traffic shapers and synchronization between TSN capable switches. These switches are connected through Ethernet and ensure connectivity between RSUs, and between RSUs and the smart city cloud.
@@ VI-B Proposal of evaluation architecture
This testbed shall include at least one or more ECUs, a data center [139], a vehicular edge computing area [139], one vehicular sensor, one vehicular actuator, a Ethernet backbone with TSN capable switch, and a CAN legacy bus.
The testbed may be deployed either by software or hardware. Software simulators can be used to mimic: a) the network, b) the vehicular topology, and c) the inter vehicular domain, while hardware can be used to function/emulate ECUs, Ethernet switches with TSN support, CAN based sensors and actuators.
Analysis of overall performance of the system should be at least in terms of throughput and latency for edge (intra-vehicular), cloud, and hybrid approaches but also consider additional metrics, such as, inter-frame/packet latency [140], data rate, transmission queue length [141], percentage of missed deadline or offset of period and worst frame/packet latency.
VI-B1 Hardware equipment
The hardware equipment that may be considered within the testbed in fig. 22 shall consider hardware equipment in two different contexts:
  • The network infrastructure, comprised by the switches, routers and access points that shall be compatible with deterministic communication standards, such as TSN standards.
  • The ITS intra vehicle hardware, including generic computers or specific computers (i.e. ECUs).
@@ VI-B1 Hardware equipment
Regarding the network infrastructure, solutions from closed source from typical network manufacturers to open source solutions are available. In principle all equipment supporting the basic architecture within [142, 143] are enough to support the basic testing of the proposed architecture. Cisco possesses TSN capable switches, such as the IE 4000 Series Industrial Switch. Within its operating system, the IOS, it is possible to install support to TSN standards [^63] Details available at https://www.cisco.com/c/en/us/td/docs/switches/lan/cisco_ie4000/tsn/b_tsn_ios_support/b_tsn_ios_support_chapter_010.html..
Closed source alternatives include for example the NXP SJA1105TEL [^64] Details at https://www.nxp.com/products/interfaces/ethernet-/automotive-ethernet-switches/five-ports-avb-and-tsn-automotive-ethernet-switch:SJA1105TEL. switch, specifically target at vehicular ECUs, or the Broadcom BCM53570 SoC.
Relyum offer closed-sourced FPGA based alternative solutions [^65] More details at https://www.relyum.com/web/rely-tsn-lab/. Equipment available at the group lab. for TSN switches, with theoretical support for configuration through NETCONF, with alternatives from SoC-e [^66] Such as https://soc-e.com/avb-automotive-ethernet-switch-ip-core/. also available.
In addition, several SoC alternatives also support TSN. Xilinx SoC, such as the KR260 series [^67] Details available at https://www.xilinx.com/content/dam/xilinx/ publications/product-briefs/kr260-product-brief.pdf. contain support for TSN, integrate a FPGA that may support the hardware acceleration of parts of the proposed architecture in fig. 19 and support the interconnection with existing legacy vehicular systems through its support for CAN bus, while also keeping an open architecture having in mind the reconfigurability FPGAs support.
@@
PixKit [^68] Details available at https://pixmoving-moveit.github.io/pixkit-documentation-en/pix-chassis-user-manual/. Equipment available on the lab as of April 2024. is a possibility currently under study. Alternatives such as the The Car Lab’s HYDRAONE [^69] Available at http://weisongshi.org/hydraone/. could also be considered, even if less advanced than the PixKit alternative, in particular without out-of-the-box semi autonomous driving [^70] Semi autonomous since the PixKit solution is fully dependent on previously mapped environments.. Within the vehicle, computation systems such as NXP, through its BlueBox Development Kit [^71] Details available at https://www.nxp.com/design/design-center/development-boards-and-designs/bluebox-3-0-automotive-high-performance-compute-ahpc-development-platform:BlueBox. or Renesas through its R-Car H3e-2G Starter Kit [^72] Details available at https://www.renesas.com/in/en/products/automotive-products/automotive-system-chips-socs/rtp8j779m1askb0sk0sa003-r-car-h3e-2g-starter-kit. can be considered as ECU for testing purposes. Well known alternatives such as the Raspberry Pi and equivalents can also be considered.
@@
Finally, one should consider hardware equipment related to the use cases themselves. Sensors and actuators that are needed, for example for the use case in section V-B) include step motors — for the ramp or fixing the seat belts — or sensors to detect unbuckled seat belts, weight sensors to detect used wheelchairs and screens for the direct interaction with the users.
VI-B2 Software and simulators
When hardware equipment is not available or if does not correspond to the requirements of the testbed as depicted, software simulation alternatives may be considered.
While solutions such as AUTOSAR simulate and model the automotive Ethernet, it lacks TSN support for such modelation [4]. To simulate the network, both in its wired and wireless components, one can use use case agnostic solutions such as OMNET [^73] Available at http://omnetpp.org/., that with INET [^74] Available at https://inet.omnetpp.org/. and with CoRE4INET [^75] Albeit with notable limitations, such as no simulation of frame preemption or clock synchronization protocols, with TSimNet model being an alternative for both frame preemption and replication and elimination [4]. or NeSTiNg [^76] Available at https://omnetpp.org/download-items/NeSTiNg.html. Most logic related to NeSTiNg is now integrated within INET. can be used to simulate networks with TTEthernet and TSN protocols [144].
OMNET can also be extended with SUMO [^77] Available at https://eclipse.dev/sumo/. through frameworks such as Veins [^78] Available at https://veins.car2x.org/., an open source framework for running vehicular network simulations [145]. It can also be extended for service-oriented communication, through the SOA4CoRE [^79] Service-Oriented Architecture for Communication over Realtime Ethernet, available at https://github.com/CoRE-RG/SOA4CoRE. initiative. Mininet, a popular alternative to create virtual networks with complex typologies and greatly used in evaluation for SDN contexts, can also be extended to considered wireless communications through the Mininet-WiFi [^80] Available athttps://mininet-wifi.github.io/. initiative and can therefore also be an alternative situation.
@@ VI-B2 Software and simulators
RealTime-At-Work Pegase [^81] Available at https://www.realtimeatwork.com/rtaw-pegase/. provides a closed proprietary tool for the equivalent purpose of testing ECUs communications. Renault, for example, considers within their TSN tests an ECU based on AUTOSAR classic software and interconnected with a physical + MAC Ethernet layer with TSN capable hardware [19]. The RTaW tool provide the Yet Another Next Generation (YANG) files for the hardware, with Bosch creating plugins to emulate detailed modeling of the ECUs integrate into the RTaW Pegase simulation [19].
@@
Tools like Autoware [^82] Available at https://autowarefoundation.github.io/autoware-documentation/main/. and Carla [^83] Available at https://carla.org. can then be used to simulate vehicular use cases, particularly in autonomous driving scenarios. Autoware in particular is considered a good solution both for being based on the widely used ROS and for offering hardware [^84] Available at https://www.autoware.org/autoware-open-ad-kit. kits that can be integrated within their software. However, while it has monitoring tools for CPU and GPU usage and some redundancy in sensor fault management, lacks any in-vehicle networking, both wired and wireless [^85] Current focus of Autoware V2X module, available at https://tlab-wide.github.io/AutowareV2X/main/..
@@
CSS Electronics CANedge2868686Details at https://www.csselectronics.com/products/can-bus-data-logger-wifi-canedge2. and equivalent equipment can be used to obtain real frames from CAN bus, either simulated or real, and send it to another computer. This can be used to obtain data to support the simulations with the aforementioned tools. If not possible, buses like CAN can also be simulated for example using Linux support for Virtual CAN tunnel (VXCAN). [^87] Source code available at https://github.com/torvalds/linux/blob/master/ drivers/net/can/vxcan.c.
VII Conclusions
In this tutorial, the concept of Software Defined Vehicles was presented and framed as essential to support the development of the next generation of vehicular services. Its main building blocks - Service Oriented Architecture, operating systems, virtualization and orchestration techniques and inter- and intra-vehicular communications with determinism techniques can support the creation of a truly Software-Defined Vehicle. An architecture was proposed for the dynamic management of services within SDV. Based on a deterministic network configurator, a data layer configurator, a hypervisor configurator and a vehicle abstraction layer, all coordinated by a software orchestrator, it supports an easier dynamic management of vehicular services, easing the development burden of such artifacts. Future work includes implementation and the proper evaluation of its scheduling algorithms as well as a end-to-end acceptance test.

### v2

@@ 6.1 Overall end-to-end interaction
Our vision includes a system that manages service deployment, execution and evaluation of intra- and inter- vehicular services. Services can be associated with domains, that define not only requirements but the universe of information the service can subscribe. Service developers define a service descriptor (an example of the current vision for its structure is detailed in listing LABEL:listing:arch-use-case-service_descriptor) to describe the main service characteristics. This set of requirements is used to configure and dispatch the services to the appropriated orchestrator in the appropriated ECUs and configure the datapath, i.e. the network and data layer resources required for the service to communicate with others.
@@
Services must be deployed by the controller which must decide the best node(s) for the service characteristics, either automatically, manually, or based on criteria (e.g. time, energy, cost). The controller must further ensure latency bounds for the deployment of the service and for the exchange of messages between them if the service descriptor requires criticality and/or QoS levels on the data exchange and processing. The controller is the module responsible for a) evaluating the nodes on the match between the service characteristics and the node characteristics, and b) dispatching the configurations to the $N$ nodes best fulfilling the service characteristics. Receiving as input both the service descriptor and the node characteristics, the algorithm outputs a score to each node and if the node may be used or not (to support requirements that must be fulfilled).
@@
The vehicular service developer shall interact with the system through a declarative paradigm, i.e. set the wanted state [^63] Kubernetes is a widely known example of such declarative paradigm.. A service descriptor file — or equivalent structure — describing the necessary characteristics of the vehicular service can support this paradigm. This information is defined by the service developer and can be used by the system to be developed to support the service deployment.
@@
Associated with each field is not only the value, with the type prescribed, but also its criticality. Since not all conditions can be met by a node, it is essential for the controller module to know which metrics can be unmet and by what margin.
@@
For example, when considering the use cases in section 5.3.2, one can imagine the service developer creating a service descriptor such as the one in listing LABEL:listing:arch-use-case-service_descriptor. This service descriptor is then submitted to the controller within the target vehicle(s). Within each vehicle a controller is responsible fo evaluating the ECUs and configuring them and the necessary network and computing components to ensure that the service requirements are properly considered. In the case of the considered service, this entails at least:
  • 1.
    Deploying the necessary containers;
  • 2.
    Allocating and configuring the necessary resources within an ECU connected to the necessary actuators (in this case the motors for the fixing points and the ramp);
  • 3.
    Configuring the switches and other network equipment connecting such actuators and the ECUs where the service connect with the necessary deterministic networking mechanisms (in this case TSN TAS scheduling could be a possibility).
while keeping fulfilling the other already running services and their requirements.
@@
6.2 Proposal of an evaluation architecture and metrics
The evaluation of the proposed system requires at least a testbed with a topology equivalent to the one depicted in fig. 24. In this topology, a vehicle form a deterministic network domain and can connect to other vehicles and smart city infrastructure. A smart city infrastructure, including its cloud and RSUs form another deterministic network domain [85]. Each domain is responsible for ensuring a deterministic end-to-end latency for messages of vehicular services by configuring appropriately the necessary traffic shapers and synchronization between TSN capable switches. These switches are connected through Ethernet and ensure connectivity between the ECUs (in the vehicle domain) or between RSUs, and between RSUs and the smart city cloud (in the smart city cloud).
@@ 6.2 Proposal of an evaluation architecture and metrics
In terms of required equipment, such a testbed shall include at least one or more ECUs, a data center [141], a vehicular edge computing area [141], one vehicular sensor, one vehicular actuator, a Ethernet backbone with TSN capable switch, and a CAN legacy bus. The testbed may be deployed either by software or hardware. Software simulators can be used to mimic: a) the network, b) the vehicular topology, and c) the inter vehicular domain, while hardware can be used to function/emulate ECUs, Ethernet switches with TSN support, CAN based sensors and actuators.
Analysis of overall performance of the system should be at least in terms of throughput and latency for edge (intra-vehicular), cloud, and hybrid approaches but also consider additional metrics, such as, inter-frame/packet latency [142], data rate, transmission queue length [143], percentage of missed deadline or offset of period and worst frame/packet latency. In terms of metrics it can include average node (ECU) pressure, uptime and downtime of each service (and associated error) and time until recovery.
6.2.1 Hardware equipment
The hardware equipment that may be considered within the testbed in fig. 24 shall consider hardware equipment in two different contexts: the network infrastructure, comprised by the switches, routers and access points that shall be compatible with deterministic communication standards, such as TSN standards; and the Intelligent Transport System (ITS) intra vehicle hardware, including generic computers or specific computers (i.e. ECUs).
@@ 6.2.1 Hardware equipment
Regarding the network infrastructure, solutions from closed source from typical network manufacturers to open source solutions are available. In principle all equipment supporting the basic architecture within [144, 145] are enough to support the basic testing of the proposed architecture. Cisco possesses TSN capable switches, such as the IE 4000 Series Industrial Switch. Within its operating system, the IOS, it is possible to install support to TSN standards [^64] Details at cisco.com..
Closed source alternatives include for example the NXP SJA1105TEL [^65] Details at nxp.com. switch, specifically target at vehicular ECUs, or the Broadcom BCM53570 System-on-chip (SoC).
Relyum offers closed-sourced FPGA based alternative solutions [^66] More details at relyum.com. for TSN switches, with theoretical support for configuration through NETCONF, with alternatives from SoC-e [^67] Such as the alternative at soc-e.com. also available.
In addition, several SoC alternatives also support TSN. Xilinx SoC, such as the KR260 series [^68] Details at xilinx.com. contain support for TSN, integrate a FPGA that may support the hardware acceleration of parts of the proposed architecture in fig. 21 and support the interconnection with existing legacy vehicular systems through its support for CAN bus, while also keeping an open architecture having in mind the reconfigurability that FPGAs support.
@@
PixKit [^69] Details at pixmoving-moveit.github.io. is a possibility currently under study. Alternatives such as the The Car Lab’s HYDRAONE [^70] Available at weisongshi.org/hydraone. could also be considered, even if less advanced than the PixKit alternative, in particular without out-of-the-box semi autonomous driving [^71] Semi autonomous since the PixKit solution is fully dependent on previously mapped environments.. Within the vehicle, computation systems such as NXP, through its BlueBox Development Kit [^72] Details at nxp.com. or Renesas through its R-Car H3e-2G Starter Kit [^73] Details at renesas.com. can be considered as ECU for testing purposes. Well known alternatives such as the Raspberry Pi and equivalents can also be considered.
@@
Finally, one should consider hardware equipment related to the use cases themselves, namely sensors and actuators. For example, the use case in section 5.3.2) includes step motors — for the ramp or fixing the seat belts — or sensors to detect unbuckled seat belts, weight sensors to detect used wheelchairs and screens for the direct interaction with the users.
6.2.2 Software and simulators
When hardware equipment is not available or if it does not correspond to the requirements of the testbed as depicted, software simulation alternatives may be considered.
While solutions such as AUTOSAR simulate and model the automotive Ethernet, they lack TSN support for such modelation [2]. To simulate the network, both in its wired and wireless components, one can use use case agnostic solutions such as OMNET [^74] At omnetpp.org., that with INET [^75] Available at inet.omnetpp.org. and with CoRE4INET [^76] Albeit with notable limitations, such as no simulation of frame preemption or clock synchronization protocols, with TSimNet model being an alternative for both frame preemption and replication and elimination [2]. or NeSTiNg [^77] Most logic related to NeSTiNg is now integrated within INET. can be used to simulate networks with TTEthernet and TSN protocols [146].
OMNET can also be extended with SUMO [^78] Available at eclipse.dev/sumo. through frameworks such as Veins [^79] At veins.car2x.org., an open source framework for running vehicular network simulations [147]. It can also be extended for service-oriented communication, through the SOA4CoRE [^80] Service-Oriented Architecture for Communication over Realtime Ethernet, at github.com/CoRE-RG/SOA4CoRE. initiative. Mininet, a popular alternative to create virtual networks with complex typologies and greatly used in evaluation for SDN contexts, can also be extended to consider wireless communications through the Mininet-WiFi [^81] At mininet-wifi.github.io. initiative and can therefore also be an alternative situation.
@@ 6.2.2 Software and simulators
RealTime-At-Work Pegase [^82] At realtimeatwork.com/rtaw-pegase. provides a closed proprietary tool for the equivalent purpose of testing ECUs communications. Renault, for example, considers within their TSN tests, an ECU based on AUTOSAR classic software and interconnected with a physical and MAC Ethernet layer with TSN capable hardware [14]. The RTaW tool provides the Yet Another Next Generation (YANG) files for the hardware, with Bosch creating plugins to emulate detailed modeling of the ECUs integrate into the RTaW Pegase simulation [14].
@@
Tools like Autoware [^83] At autowarefoundation.github.io. and Carla [^84] At carla.org. can then be used to simulate vehicular use cases, particularly in autonomous driving scenarios. Autoware in particular is considered a good solution both for being based on the widely used ROS and for offering hardware [^85] Available at autoware.org/autoware-open-ad-kit. kits that can be integrated within their software. However, while it has monitoring tools for CPU and GPU usage and some redundancy in sensor fault management, it lacks any in-vehicle networking, both wired and wireless [^86] Current focus of Autoware V2X module, details at tlab-wide.github.io/AutowareV2X..
@@
CSS Electronics CANedge2878787Details at csselectronics.com. and equivalent equipment can be used to obtain real frames from CAN bus, either simulated or real, and send it to another computer. This can be used to obtain data to support the simulations with the aforementioned tools. If not possible, buses like CAN can also be simulated for example using Linux support for Virtual CAN tunnel (VXCAN). [^88] Source code available at github.com/torvalds/linux/blob/master/ drivers/net/can/vxcan.c..
7 Conclusions
In this article, the concept of Software Defined Vehicles was presented and framed as essential to support the development of the next generation of vehicular services. Considering the criticality of the future SDV services, as well as the increased complexity of development and deployment of services, it is more and more critical to consider a dynamic management of services.
The main building blocks of SDV to be considered in such a dynamic management were listed and detailed, as well as the main associated challenges. Vehicles will tend to a physical zonal architecture, supporting a set of services running in a set of higher performance ECUs (in a SOA paradigm). Existing operating systems will have to evolve towards interoperability and supporting virtualization and orchestration techniques, as well as integrating inter- and intra-vehicular communications with determinism techniques to support critical services that can have direct impact of the occupants and surroundings of the vehicles.
While the industry, including automakers and their suppliers, are working towards SDVs, their initiatives are still tending to be proprietary or closed but the trend is towards greater interoperability with holistic solutions such as ”the collaboration of collaborations” between AUTOSAR, COVESA, Eclipse SDV and SOAFEE announced in CES 2024.
A vision towards the dynamic management of services within SDV. Based on a deterministic network configurator, a data layer configurator, a orchestrator and a vehicle abstraction layer, all coordinated by a controller, it supports an easier dynamic management of vehicular services, easing the development burden of such artifacts.
Future work towards such vision includes the further specification, service orchestration, implementation and the proper evaluation of its scheduling algorithms, as well as an end-to-end acceptance tests.